OpenStudy (anonymous):

Stupid question, but oh well. This is what we're supposed to do: Write each answer as a power of 2. Show your work and explain your steps. O.K, I got the answer of 131072, but I don't know how to make that a power of 2. My work: 2^20/2^3. How do you turn 131,072 into a power of 2? T-T

OpenStudy (anonymous):

\[131072\] is an fact a power of 2 (who knew?)

OpenStudy (anonymous):

keep dividing by 2 until you get it is one way i would cheat

OpenStudy (anonymous):

actually, you already have it, but your answer is more convoluted than necessary since \[\frac{2^{20}}{2^3}=2^{20-3}=2^{17}\]

OpenStudy (anonymous):

my question is, how in the world did you find it was \[\frac{2^{20}}{2^3}\]??

OpenStudy (anonymous):

Oh I see DX Dang it, I didn't even think of the same base thing. I got the 2^30/2^3 because it was in the problem, and the megabytes and bits and just ugh. Lol. Thanks for helping me though.

OpenStudy (anonymous):



OpenStudy (anonymous):

Whoa that's helpful. Thanks!